# Phase reduction explains chimera shape: when multi-body interaction matters

### v2

@@ Abstract
We present an extension of the Kuramoto-Sakaguchi model for networks, deriving the second-order phase approximation for a paradigmatic model of oscillatory networks - an ensemble of non-identical Stuart-Landau oscillators coupled pairwisely via an arbitrary adjacency matrix. We explicitly demonstrate how this matrix translates into the coupling structure in the phase equations. To illustrate the power of our approach and the crucial importance of high-order phase reduction, we tackle a trendy setup of non-locally coupled oscillators exhibiting a chimera state. We reveal that our second-order phase model reproduces the dependence of the chimera shape on the coupling strength that is not captured by the typically used first-order Kuramoto-like model. Our derivation contributes to a better understanding of complex networks’ dynamics, establishing a relation between the adjacency matrix and multi-body interaction terms in the high-order phase model.
@@ II Main results
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ III Second-order phase reduction for the SL network
For convenience, we first switch to physically meaningful parameters for each unit: frequency of the limit-cycle oscillation, $\omega_{i}=\nu_{i}-\eta_{i}\beta_{i}$, Floquet exponent of the cycle, $\kappa_{i}=-2\eta_{i}<0$, and the non-isochronicity parameter $\gamma_{i}=\arctan(\beta_{i})$.
With $c_{i}=\kappa_{i}\tan(\gamma_{i})/2$, Eq. (1) becomes
@@
with $\tan(\gamma)=2c/\kappa$, which is defined by two characteristic properties. First, in the new coordinates, the dynamics of the amplitude $r$ and the phase $\phi$ decouple and take on a simple linear form $\dot{\phi}=\omega$ and $\dot{r}=\kappa r$. Second, on the limit cycle ($\rho=1$), it holds $r=0$ and $\phi=\theta$. Note that the transformation (9,10) defined in this way  [^3] The phase-isostable transformation (9,10) is obtained by solving two differential equations $\dot{\phi}=\omega=\frac{d}{dt}\Phi(\theta,\rho)$ and $\dot{r}=\kappa r=\frac{d}{dt}P(\rho)$, which explicitly read $\omega=(\omega+c(\rho^{2}-1))\partial_{\theta}\Phi+0.5\kappa\rho(\rho^{2}-1)\partial_{\rho}\Phi$ and $\kappa P=0.5\kappa\rho(\rho^{2}-1)\partial_{\rho}P$, with two boundary conditions $\Phi(\theta,1)=\phi$ and $P(1)=0$. Note that although Eq. (9) is a unique solution, Eq. (10) is defined only up to a scaling factor. is isomorphic on the entire basin of attraction of the limit cycle, i.e. for all $\rho>0$ [25]. In addition, the isochrones of the SL oscillator $\Phi(\theta,\rho)=\mathrm{const}$ are logarithmic spirals with the inclination angle $\gamma$ relative to the outward normal on the limit cycle ($\rho=1$). For $\gamma=0$, these spirals turn into straight rays and Eq. (9) degenerates into the identity $\phi=\theta$.
@@ III.1 A particular case: global coupling
-- $\cos(\Delta_{jk})\sin(2\phi_{k}-\phi_{j}-\phi_{i}+\Delta_{jk})$
-- $\cos(\Delta_{jk})\sin(\phi_{j}-\phi_{i}-\Delta_{jk}+2\alpha)\biggr{)}\,.$ (24)
@@ IV Discussion
We summarize our results. We presented second-order phase reduction for a network of non-identical and arbitrarily coupled SL oscillators. The main effect is that pairwise coupling of the SL units results in the emergence of hypernetwork with triplet interactions on the level of phase description. Our derivation yields the correct structure of multi-body interaction in the physically motivated extension of the Kuramoto-Sakaguchi model on networks, in particular, for globally coupled units, see Eq. (24). Furthermore, we provide the phase equation describing chimera states with an account of second-order terms. The relation of the coupling strength and the Floquet multiplier of the cycle determines the amplitude of these terms. Our derivation implies that the second-order terms are small compared to the KS terms only for $\varepsilon\ll|\kappa|$, while they cannot be neglected for moderate coupling. Generally, high-order phase reduction yields qualitative and quantitative improvement in describing the dynamics. The illustration for the former case is increased precision in determining the synchronization domain of the harmonically forced van der Pol oscillator  [24]. Examples for the latter case include an explanation of the remote synchrony in a motif of three coupled SL systems [32] and coupling-dependent effects in chimera dynamics shown in Fig. 1.
Finally, we mention that general results [24] provide a route to computing third- and higher-order terms, leading to quadruplet and so on terms, though this task is tedious and highly time-consuming. Another conclusion is that multi-body interaction appears for networks of arbitrary limit-cycle oscillators, though the closed analytical derivation of the second-order terms becomes unfeasible and must be partially done numerically, cf. [24]. We emphasize that there exists an entirely different scenario resulting in multi-body interaction terms due to nonlinear coupling [33, *Komarov-Pikovsky-15a, *ashwin2016, *Pikovsky-Rosenblum-22] which we do not address here.

### v4

@@ Abstract
We present an extension of the Kuramoto-Sakaguchi model for networks, deriving the second-order phase approximation for a paradigmatic model of oscillatory networks - an ensemble of non-identical Stuart-Landau oscillators coupled pairwisely via an arbitrary coupling matrix. We explicitly demonstrate how this matrix translates into the coupling structure in the phase equations. To illustrate the power of our approach and the crucial importance of high-order phase reduction, we tackle a trendy setup of non-locally coupled oscillators exhibiting a chimera state. We reveal that our second-order phase model reproduces the dependence of the chimera shape on the coupling strength that is not captured by the typically used first-order Kuramoto-like model. Our derivation contributes to a better understanding of complex networks’ dynamics, establishing a relation between the coupling matrix and multi-body interaction terms in the high-order phase model.
@@ III Second-order phase reduction for the SL network
For convenience, we first switch to physically meaningful parameters for each unit: frequency of the limit-cycle oscillation, $\omega_{i}=\nu_{i}-\eta_{i}\beta_{i}$, Floquet exponent of the cycle, $\kappa_{i}=-2\eta_{i}<0$, and the non-isochronicity parameter $\gamma_{i}=\arctan(\beta_{i})$. With $c_{i}=\kappa_{i}\tan(\gamma_{i})/2$, Eq. (1) becomes
@@
with $\tan(\gamma)=2c/\kappa$, which is defined by two characteristic properties. First, in the new coordinates, the dynamics of the amplitude $r$ and the phase $\phi$ decouple and take on a simple linear form $\dot{\phi}=\omega$ and $\dot{r}=\kappa r$. Second, on the limit cycle ($\rho=1$), it holds $r=0$ and $\phi=\theta$. Note that the transformation (9,10) defined in this way  [^3] The phase-isostable transformation (9,10) is obtained by solving two differential equations $\dot{\phi}=\omega=\frac{d}{dt}\Phi(\theta,\rho)$ and $\dot{r}=\kappa r=\frac{d}{dt}P(\rho)$, which explicitly read $\omega=(\omega+c(\rho^{2}-1))\partial_{\theta}\Phi+0.5\kappa\rho(\rho^{2}-1)\partial_{\rho}\Phi$ and $\kappa P=0.5\kappa\rho(\rho^{2}-1)\partial_{\rho}P$, with two boundary conditions $\Phi(\theta,1)=\theta$ and $P(1)=0$. Note that although Eq. (9) is a unique solution, Eq. (10) is defined only up to a scaling factor. is isomorphic on the entire basin of attraction of the limit cycle, i.e. for all $\rho>0$ [25]. In addition, the isochrones of the SL oscillator $\Phi(\theta,\rho)=\mathrm{const}$ are logarithmic spirals with the inclination angle $\gamma$ relative to the outward normal on the limit cycle ($\rho=1$). For $\gamma=0$, these spirals turn into straight rays and Eq. (9) degenerates into the identity $\phi=\theta$.
@@ III.1 A particular case: global coupling
-- $\cos(\Delta_{jk})\sin(2\phi_{j}-\phi_{k}-\phi_{i}+\Delta_{jk})$
-- $\cos(\Delta_{jk})\sin(\phi_{k}-\phi_{i}-\Delta_{jk}+2\alpha)\biggr{)}\,.$ (24)
@@ IV Discussion
We summarize our results. We presented second-order phase reduction for a network of non-identical and arbitrarily coupled SL oscillators. The main effect is that pairwise coupling of the SL units results in the emergence of hypernetwork with triplet interactions on the level of phase description. Our derivation yields the correct structure of multi-body interaction in the physically motivated extension of the Kuramoto-Sakaguchi model on networks, in particular, for globally coupled units, see Eq. (24). Furthermore, we provide the phase equation describing chimera states with an account of second-order terms. The relation of the coupling strength and the Floquet multiplier of the cycle determines the amplitude of these terms. Our derivation implies that the second-order terms are small compared to the KS terms only for $\varepsilon\ll|\kappa|$, while they cannot be neglected for moderate coupling. Generally, high-order phase reduction yields qualitative and quantitative improvement in describing the dynamics. The illustration for the former case is increased precision in determining the synchronization domain of the harmonically forced van der Pol oscillator  [24]. Examples for the latter case include an explanation of the remote synchrony in a motif of three coupled SL systems [32], coupling-dependent effects in chimera dynamics shown in Fig. 1, and delay-induced synchronization [33].
In addition, we emphasize that the demonstrated effect of high-order terms on the chimera’s shape applies to a much broader range of models and phenomena. Indeed, the standard theory says that for general limit-cycle oscillators and arbitrary network topology, the first-order approximation (the Kuramoto-Daido form) reads $\dot{\phi}_{i}=\omega_{i}+\varepsilon\mathcal{R}_{i}(\vec{\phi}-\phi_{i}\vec{e})$, where $\mathcal{R}_{i}(\vec{\phi}-\phi_{i}\vec{e})$ are $2\pi$-periodic functions which generally contain high harmonics, not only the first one. Nonetheless, in the case of identical oscillators $\omega_{i}=\omega$, we again eliminate the coupling parameter $\varepsilon$ by moving to the co-rotating frame and rescaling time. Therefore, a Kuramoto-Daido first-order phase model cannot, in principle, describe the effect of the coupling strength on the shape of any coherence-incoherence pattern in a network. Thus, incorporating the second-order terms $\sim\varepsilon^{2}$ becomes crucial.
Finally, we mention that general results [24] provide a route to computing third- and higher-order terms, leading to quadruplet and so on terms, though this task is tedious and highly time-consuming. Another conclusion is that multi-body interaction appears for networks of arbitrary limit-cycle oscillators, though the closed analytical derivation of the second-order terms becomes unfeasible and must be partially done numerically, cf. [24]. We emphasize that there exists an entirely different scenario resulting in multi-body interaction terms due to nonlinear coupling [34, *Komarov-Pikovsky-15a, *ashwin2016, *Pikovsky-Rosenblum-22] which we do not address here.